a psychologist has designed a questionnaire to measure individuals aggressiveness. Suppose at the scores on the questionnaire or normally distributed with the standard deviation of 90 suppose also that exactly 10% of the scores exceed 750. Find the mean of the distribution of scores care your intermediate complications to at least four decimal places. Round your answer is to at least one decimal place
Answer #1

Solution:

P(Z < (750-mu)/90) = 0.1
(750-mu)/90) = 1.28
750 = 1.28*90
750-mu = 115.2
-mu = 115.2-750
mu = 750-115.2
mu = 634.8

=>  the mean of the distribution of scores is 634.8